# Making Them a Malicious Database: Exploiting Query Code to Jailbreak Aligned Large Language Models

### v2

@@ 4.4 Discussion about Countermeasures
We consider two baseline defenses and design a tailored defense method against QueryAttack as follows. Detailed settings for these defense methods can be found in the Appendix B.2.
@@
Cross-lingual Alignment Prompting based defense (Ours):  (Qin et al., 2023) propose a method that uses cross-lingual chain-of-thought (CoT) prompting to generate reasoning paths, improving zero-shot CoT reasoning across languages. Although this approach is not originally designed for defense, we find that such CoT reasoning can help LLMs recognize cross-lingual malicious intent. Our insight is that, the success of QueryAttack relies on the target model’s ability to accurately interpret custom language templates, meaning the model should also be capable of identifying the intent of them and translating them into natural language. Once translated into natural language, malicious prompts are more likely to be filtered by existing safety alignment defenses. This indicates that a well-executed translate-then-reason CoT process can effectively defend against QueryAttack-like jailbreak attacks. Based on this, we design a defense method using cross-lingual chain-of-thought prompting. The complete chain of thought is: we first requires the target model to identify the Query Content, Key Object, and Content Source of the input, then describe the query in natural language. The target model then responds to this natural language query, thereby activating existing safety alignment mechanisms for defense. The detailed prompt can be found in the Appendix B.2.
@@ 6 Conclusion
In this paper, we investigate the generalization challenges faced by large language models with safety alignment when encountering out-of-distribution malicious structured non-natural query language. Specifically, we introduce QueryAttack, a novel jailbreak attack framework. QueryAttack extracts three query components from a query in natural language, fill them into query templates of various styles, and leverages the obtained query code to bypass the target LLM’s safety alignment. Although QueryAttack does not encrypt the outputs, our extensive evaluation shows that it still effectively bypasses the defenses of mainstream LLMs and can withstand common defense methods. Besides, to defend against QueryAttack, we develop a tailored defense strategy. Experimental results demonstrate that it effectively enhances the ability of mainstream models to resist such attacks.
@@
Appendix A Attack Examples of QueryAttack
A.1 Success Examples
Figure 6: An attack example of QueryAttack using a C++ style template.
A.2 Prompts of Query Learning
@@
Appendix B Experimental Details
B.1 Human Evaluation on GPT-4 Judge
@@ B.2 Defense Settings
B.3 Ethic-related Disclaimer Removal
@@
Appendix C Prompts for Extracting Query Components

### v3

@@ 4.4 Discussion about Countermeasures
We consider two baseline defenses and design a tailored defense method against QueryAttack as follows. Detailed settings for these defense methods can be found in the Appendix C.2.
@@
Cross-lingual Alignment Prompting based defense (Ours):  (Qin et al., 2023) propose a method that uses cross-lingual chain-of-thought (CoT) prompting to generate reasoning paths, improving zero-shot CoT reasoning across languages. Although this approach is not originally designed for defense, we find that such CoT reasoning can help LLMs recognize cross-lingual malicious intent. Our insight is that, the success of QueryAttack relies on the target model’s ability to accurately interpret custom language templates, meaning the model should also be capable of identifying the intent of them and translating them into natural language. Once translated into natural language, malicious prompts are more likely to be filtered by existing safety alignment defenses. This indicates that a well-executed translate-then-reason CoT process can effectively defend against QueryAttack-like jailbreak attacks. Based on this, we design a defense method using cross-lingual chain-of-thought prompting. The complete chain of thought is: we first requires the target model to identify the Query Content, Key Object, and Content Source of the input, then describe the query in natural language. The target model then responds to this natural language query, thereby activating existing safety alignment mechanisms for defense. The detailed prompt can be found in the Appendix C.2.
@@ 6 Conclusion
In this paper, we investigate the generalization challenges faced by large language models with safety alignment when encountering out-of-distribution malicious structured non-natural query language. Specifically, we introduce QueryAttack, a novel jailbreak attack framework. QueryAttack extracts three query components from a query in natural language, fill them into query templates of various styles, and leverages the obtained query code to bypass the target LLM’s safety alignment. Although QueryAttack does not encrypt the outputs, our extensive evaluation shows that it still effectively bypasses the defenses of mainstream LLMs and can withstand common defense methods.
Besides, to defend against QueryAttack, we develop a tailored defense strategy. Experimental results demonstrate that it effectively enhances the ability of mainstream models to resist such attacks.
Acknowledgments
We thank the anonymous reviewers for their feedback on this work. This work is supported by the Major Key Project of PCL under grant No. PCL2023A06, Shenzhen R&D Program under grant NO. KJZD20230923114059020, and in part by Shenzhen Key Laboratory of Software Defined Networking under Grant ZDSYS20140509172959989.
@@ Acknowledgments
Appendix A Evaluation about the Capability of LLMs in Understanding and Generating Encrypted Content
To evaluate the capability of LLMs in understanding and generating encrypted content, we adopt the prompt used in CipherChat (Yuan et al., 2024) for Caesar ciphers and construct a benign request (as malicious ones might trigger defense mechanisms): “Please tell me a story for children. This story should include a big bad wolf and a little girl.” We then translate it into a Caesar cipher with a shift of 3 and submitted it to mainstream LLMs. We evaluate the models’ responses from two perspectives: (1) whether the model could accurately understand the encrypted text, and (2) whether it could generate a high-quality response that meets user expectations (i.e., free of readability issues and logically coherent). The results are shown in Table 5.
Model GPT-4o Gemini-1.5-Pro Llama-3.1-8B DeepSeek-R1 GPT-4.5
Understanding
Correctness
Yes No No Yes Yes
Generation
Quality
No No No No Yes
Table 5: Capability of LLMs in Understanding and Generating Encrypted Text.
As an example, here is a response from GPT-4o (after translated into natural language): “It looks like a fair| tale! Here’s a nice mi[ of a big bad wolf and a little girl: 1. Big bad wolf = This could be a reference to a truck, an anquatic and dangerods one. 2. Little girl = Ma|be a little girl is a compact temple and a late gravitation surge. You might be pointing to a scene from a sci-fi movie or book—specificall| a polot involving a large and dangerods poawer fuell incorporated into a temple with a compact surge s
stem. Would you like me to e[plore a narrative based on this ciffer”
Appendix B Attack Examples of QueryAttack
B.1 Prompts of Query Learning
@@
B.2 Success Examples
Figure 9: An attack example of QueryAttack using a C++ style template.
Appendix C Experimental Details
C.1 Human Evaluation on GPT-4 Judge
@@ C.2 Defense Settings
C.3 Ethic-related Disclaimer Removal
@@
Appendix D Prompts for Extracting Query Components